# Modeling the dynamics of an oil drop driven by a surface acoustic wave in the underlying substrate

### v3

@@ Abstract
We present a theoretical study, supported by simulations and simple experiments, on the spreading of a silicone oil drop under MHz-frequency surface acoustic wave (SAW) excitation in the underlying solid substrate. Our time-dependent theoretical model uses the long wave approach and considers interactions between fluid dynamics and acoustic driving. While similar methods have analyzed micron-scale oil and water film dynamics under SAW excitation, acoustic forcing was linked to boundary layer flow, specifically Schlichting and Rayleigh streaming, and acoustic radiation pressure. For the macroscopic drops in this study, acoustic forcing arises from Reynolds stress variations in the liquid due to changes in the intensity of the acoustic field leaking from the SAW beneath the drop and the viscous dissipation of the leaked wave. Contributions from Schlichting and Rayleigh streaming are negligible in this case. Both experiments and simulations show that after an initial phase where the oil drop deforms to accommodate acoustic stress, it accelerates, achieving nearly constant speed over time, leaving a thin wetting layer. Our model indicates that the steady speed of the drop results from the quasi-steady shape of its body. The drop speed depends on drop size and SAW intensity. Its steady shape and speed are further clarified by a simplified traveling wave-type model that highlights various physical effects. Although the agreement between experiment and theory on drop speed is qualitative, the results’ trend regarding SAW amplitude variations suggests that the model realistically incorporates the primary physical effects driving drop dynamics.
@@ 1 Introduction
A separate mechanism, the Eckart streaming, appears regardless of the presence of a boundary. This is an acoustic streaming that occurs in the bulk fluid, and results from variations in the intensity of sound or ultrasound waves in the fluid. In the classic work by Eckart and in many subsequent studies, acoustic streaming is ascribed to the viscous attenuation of traveling waves in the fluid (Eckart, 1948; Nyborg, 1952; Lighthill, 1978). This attenuation results in spatial variations of the wave intensity in the fluid, and in corresponding spatial variations in convective Reynolds stresses. The consequence is a secondary flow, whose steady component at long times is the acoustic streaming, characterized by an intense vortical flow field that appears in the bulk fluid. This is the main acoustic mechanism for the actuation of fluid in micro-channels (Wixforth, 2003; Wixforth et al., 2004; Yeo & Friend, 2014) and in drop microfluidics (Guttenberg et al., 2004; Brunet et al., 2007, 2009, 2010a). It usually dominates contributions to the flow in the bulk fluid arising from the boundary layer type acoustic streaming mechanisms, such as Shlichting or Rayleigh streaming described above. An additional effect related to the interaction of acoustic waves with a surface, the acoustic radiation pressure (Hertz & Mende, 1939; Hamilton & Blackstock, 1998), may be also relevant in the presence of interfaces, such as the fluid-air surface. This effect has been discussed extensively in the literature (King, 1934; Campbell, 1970; Chu & Apfel, 1982; Shiokawa et al., 1989; Hasegawa et al., 2000; Borgnis, 2010; Karlsen et al., 2016; Rajendran et al., 2023), and shown capable of deforming and displacing soft interfaces and of governing the dynamics of films (Biwersi et al., 2000; Alzuaga et al., 2005; Issenmann et al., 2006; Rajendran et al., 2022).
@@
While the literature discussing the interaction of acoustic waves with fluids is clearly extensive, in what follows we discuss briefly just a few works that are particularly relevant in the present context, in which a SAW traveling in a solid substrate drives flow in a neighboring fluid. In 1970 Campbell & Jones (1970) simulated a SAW in a piezoelectric solid in contact with an ideal (inviscid) fluid. Then, in the late 1980s, Shiokawa et al. (1989) used the ansatz of a harmonic acoustic waveform in the solid based on ideas developed by Nyborg (1952), converted the approach adopted by Campbell & Jones (1970) to analytic expressions, and calculated the resulting acoustic forcing in the adjacent inviscid fluid due to attenuation of SAW in the solid. More recently, Vanneste & Bühler (2011) revisited the work of Campbell & Jones (1970) in the context of viscous fluids and calculated numerically the corresponding acoustic forcing in the fluid.
@@
In the present paper, we focus on developing a model that can be used to formulate a long wave theory applicable to the free surface evolution of thin viscous films and drops due to the presence of a SAW propagating in an underlying solid substrate. To facilitate such a development, we simplify the previously considered approaches in order to isolate the dominant physical effects governing the dynamics of the film. In particular, we simplify the approach used by Campbell & Jones (1970) by assuming damped harmonic waves in the solid. Furthermore, we compare the previous approaches by Vanneste & Bühler (2011) and Shiokawa et al. (1989) to quantify the importance of viscous effects, which were included by the former authors but ignored by the latter. Moreover, we follow earlier attempts (Brunet et al., 2010b) and focus on the contribution of acoustic streaming. In principle one could also consider including the effects of the acoustic radiation pressure at the free surface, which may deform the drop. However, since the main contribution to drop displacement is the acoustic streaming, which efficiently transfers power from the SAW to translational motion of the drop, we do not include acoustic radiation pressure in the present work.
@@
The rest of this paper is structured as follows. As a motivation, in Sec. 2 we discuss simple physical experiments that demonstrate the well-known phenomenon of thick oil films and drops displacing under the action of MHz-frequency Rayleigh-type SAWs. These experiments illustrate the process of dynamic wetting and drop displacement powered by bulk acoustic streaming. We then derive the theoretical model that can explain such observations in Sec. 3. Our theoretical results are presented in Sec. 4: we discuss computational results in Sec. 4.1, before analyzing a simplified traveling–wave type model in Sec. 4.2. Section 5 is devoted to the summary and conclusions. The appendices include (A) details of the theoretical results, (B) various simplifications and approximations including the inviscid limit, and (C) the outline of the numerical approach implemented.
@@ 2 Experiment
The experiments described here are illustrative, and portray the well-known phenomenon in which MHz-frequency surface acoustic waves (SAWs) are used to displace drops and thick films of liquid along the path of the SAW. The purpose of these experiments is to highlight the system under investigation and to provide results that will be used to verify the theory and simulations presented in the following sections.
Figure 1 shows the experimental setup (A) and a typical experimental result (a-e). We generate a propagating 20 MHz frequency SAW by applying a same-frequency sinusoidal voltage signal to a piezoelectric actuator – a SAW device. The actuator comprises a 5 nm titanium/1 $\mu$m aluminum interdigitated transducer (IDT; from which the SAW emanates) fabricated atop lithium niobate (LiNbO3, Roditi International, UK) by standard lift-off photolithography. The substrate used for the SAW device is 11 mm $\times$ 24 mm in size, 0.5 mm thick, 128° Y-cut, X-propagating, single-crystal piezoelectric lithium niobate, where X and Y are crystal axes (Campbell & Jones, 1968). The actuator is integrated into the external electrical signal using pogo pins (BC201403AD, Interconnect Devices, Inc.) assembled in a 3D-printed elastomeric stage, which holds the actuator and is connected to a signal generator (R&S SMB100A microwave signal generator) and amplifier (model A10160, Tabor Electronics Ltd.). Prior to the experiment we clean the SAW device using four different solvents: acetone (AR-b, 99.8%, 67-641, Bio-Lab Ltd.), 2-propanol (AR-b, 99.8%, 67-63-0, Bio-Lab Ltd.), ethanol (CP-p, 96%, 64-17-5, Bio-Lab Ltd.), and water (HPLC plus, 7732-18-5, Sigma-Aldrich), before drying the actuator using compressed air. We place a paper cylinder soaked with glycerol at the far end of the actuator from the IDT to absorb the SAW and prevent reflections.
Figure 1: (A) Upper schematic view of the experimental setup: Surface acoustic wave (SAW) propagates from the interdigitated transducer (IDT) until it reaches the acoustic absorber (comprised of glycerol-soaked paper placed on the actuator under and to the right of the needle). The needle is of known diameter (510 $\mu$m), placed for identifying spatial resolution in the images. (a-e) Successive snapshots are taken from an experiment monitoring the flow of a silicone oil film. During the experiment, a drop of silicone oil is placed on the horizontal surface (a), it deforms due to the application of SAW (b), and moves in the direction of SAW propagation (c-e); time is shown in seconds, and the vertical lines serve as a reference.
@@
In our experiments, we represent the SAW strength (intensity) by measuring the normal displacement amplitude, $A_{n}$, at the solid surface before any attenuation in the liquid, so that $A_{n}$ is the amplitude of the transverse component of the SAW in the substrate. We measure $A_{n}$ over a surface of $1\times 1~{}\text{mm}^{2}$ about $2$ mm away from the SAW actuator using a scanning laser Doppler vibrometer (MSA-500, Polytech). We control this amplitude by changing the applied voltage, $V$, of the signal generator at the surface of the actuator, consistent with the literature (Ballantine et al., 1996), there is a linear relationship between $A_{n}$ and $V$; see Li et al. (2025) for more details.
We place an 8 mm3 ($8~{}\mu$l) drop of silicone oil (50 cSt, 378356, Sigma-Aldrich) atop the actuator using a pipette, approximately 3.5 mm away from the IDT, and introduce electrical signal at different voltage levels to induce motion of the oil film, which dynamically wets the solid substrate along the path of the SAW. Figure 2(a-c) shows a time-lapse (side view) of the typical silicone oil dynamics, captured using the framework and camera of a goniometer (Data Physics; OCA 15Pro). Figure 2(d) shows a top view obtained using a camera (EOS R5, Canon) with a macro lens (RF 100 mm F2.8L MACRO IS USM, Canon).
@@
Figure 2: Two sets of images from experiments monitoring silicone oil drops powered by SAWs of three different measured amplitudes $A_{n}$ (1.2 nm–1.4 nm), where (A-C) silicone oil starts from the same location, moves in the direction of the SAW, and reaches different distances (a-c) after 4.09 seconds. The corresponding side-view videos at different acoustic power levels are used to analyze the speed and profile of silicone oil during the movement. (D, d) show top view snapshots (corresponding to (C, c)) of the oil. A thin film of oil, hardly visible in the side view, can be seen more clearly behind the main body of silicone oil in the top views.
Figure 3: Time evolution of (a) the front position and (b) the maximum drop height for several values of $A_{n}$. The symbols correspond to the measurements and the lines are simple fits to guide the eye. We expect errors of ±$\pm 50~{}\mu$m in the values of the film height due to the limited resolution of the side view image and tracker software.
@@
In the experiments with SAW actuation we employ the side view video, recorded by a goniometer camera, to capture the position of the advancing three-phase contact line of the SAW-actuated oil film using the open-source software Tracker (Brown & Cox, 2009) (see Fig. 3(a) for typical data). We also measure the maximum height of the moving oil film using the public domain image processing software ImageJ (typical results in Fig. 3(b)). Figure 3(a) shows that, as expected, the larger the SAW amplitude $A_{n}$, the faster the oil film moves due to the stronger induced leakage wave (Shiokawa et al., 1989). Another notable aspect of SAW-induced dynamics is the contact line speed for early times ($<1$ s): initially, it is rather small, but later increases and then remains constant. This change of contact line speed appears to be a consequence of the initial change of shape of the drop (see also Fig. 1); while this is taking place, contact line speed is small. After the initial deformation, the drop shape remains unchanged, translating approximately uniformly. Figure 3(b) shows consistent results for the maximum drop height, with an increase of the height for early times, and constant values for longer times. Note that this uniform translation and constant film height differ significantly from the observed behavior for spreading under a body force such as gravity. These experimental observations will be explained and discussed within the framework of our theoretical model, which is described next.
@@ 3 Model
Let us consider a two-dimensional Newtonian fluid layer occupying a domain in the $(x,z)$–plane, traversed by a sound wave generated by a Rayleigh-type-SAW traveling in a neighboring solid; see Fig. 4 for a sketch.
Figure 4: Schematic of a liquid drop (blue curve) driven by a leaky SAW of amplitude $A_{n}$. The SAW travels from left to right in the solid substrate. The drop extends from the rear to the front contact lines at $x_{\rm r}(t)$ and $x_{\rm f}(t)$, respectively. The thin red line represents the amplitude $A_{n}$ of the SAW at the substrate and shows how it is attenuated due to the presence of the liquid (the black dotted line is the envelope of the decaying amplitude in the liquid). The Rayleigh angle is denoted by $\theta_{R}$. The thickness $h^{\ast}$ and the positions $x^{\ast}_{1,2}$ are defined later in the text, see Eq. (34).
We assume that the associated acoustic field Mach number, $M_{a}\equiv\omega A_{n}/c$ ($\omega$, $A_{n}$ and $c$ are the SAW angular frequency, characteristic normal displacement amplitude at the solid surface, and the phase velocity of the acoustic field in the fluid, respectively) is small ($M_{a}\ll 1$), and that the velocity ($\mathbf{v}$), pressure ($p$), and density ($\rho$) fields in the fluid can be written as
@@
where the leading order components of these fields are associated with quiescent liquid ($\mathbf{v}_{0}=0$, $p_{0}$ constant) of constant ambient density $\rho_{0}$; the first-order corrections ($\mathbf{v}_{1},~{}p_{1},~{}\rho_{1}$), represent an oscillatory flow associated with the acoustic field generated by the SAW that average to zero over the period of the acoustic forcing; and the second-order corrections ($\mathbf{v}_{2},~{}p_{2},~{}\rho_{2}$) represent the induced steady streaming flow. The assumption implicit in the above expansions in the Mach number is that the barred quantities are comparable in size to the corresponding leading-order contributions. The problem is governed by the Navier–Stokes (NS) equations and the continuity equation,
@@
where $t$ is time; $\mu$, $\mu_{b}$ are the shear and dilatational coefficients of the bulk viscosity, respectively, and $\mathbf{g}=-g\mathbf{e}_{z}$ is the gravitational acceleration. In terms of the above expansions, we will assume that the gravitational term does not appear until second-order in $M_{a}$, which amounts to an assumption that $\mu v_{1}/\ell^{2}\gg\rho g$ (where $v_{1}$ is a representative value of $|\mathbf{v}_{1}|$ and $\ell$ is a representative droplet size), ensuring that gravity is negligible in both leading-order and first-order problems. This condition was checked a posteriori and found to hold. With these assumptions, we substitute the expansions from Eq. (3) into Eq. (2) and collect terms of like orders in $M_{a}$; Secs. 3.1 and 3.2 discuss the first and second order terms, respectively.

### v4

@@ Abstract
We present a theoretical study, supported by simulations and experiments, on the spreading of a silicone oil drop under MHz-frequency surface acoustic wave (SAW) excitation in the underlying solid substrate. Our time-dependent theoretical model uses the long wave approach and considers interactions between fluid dynamics and acoustic driving. While similar methods have analyzed micron-scale oil and water film dynamics under SAW excitation, acoustic forcing was linked to boundary layer flow, specifically Schlichting and Rayleigh streaming, and acoustic radiation pressure. For the macroscopic drops in this study, acoustic forcing arises from Reynolds stress variations in the liquid due to changes in the intensity of the acoustic field leaking from the SAW beneath the drop and the viscous dissipation of the leaked wave. Contributions from Schlichting and Rayleigh streaming are negligible in this case. Both experiments and simulations show that after an initial phase where the oil drop deforms to accommodate acoustic stress, it accelerates, achieving nearly constant speed over time, leaving a thin wetting layer. Our model indicates that the steady speed of the drop results from the quasi-steady shape of its body. The drop speed depends on drop size and SAW intensity. Its steady shape and speed are further clarified by a simplified traveling wave-type model that highlights various physical effects. Although the agreement between experiment and theory on drop speed is qualitative, the results’ trend regarding SAW amplitude variations suggests that the model realistically incorporates the primary physical effects driving drop dynamics.
@@ 1 Introduction
A separate mechanism, the Eckart streaming, appears regardless of the presence of a boundary. This is an acoustic streaming that occurs in the bulk fluid, and results from variations in the intensity of sound or ultrasound waves in the fluid. In the classic work by Eckart and in many subsequent studies, acoustic streaming is ascribed to the viscous attenuation of traveling waves in the fluid (Eckart, 1948; Nyborg, 1952; Lighthill, 1978). This attenuation results in spatial variations of the wave intensity in the fluid, and in corresponding spatial variations in convective Reynolds stresses. The consequence is a secondary flow, whose steady component at long times is the acoustic streaming, characterized by an intense vortical flow field that appears in the bulk fluid. This is the main acoustic mechanism for the actuation of fluid in micro-channels (Wixforth, 2003; Wixforth et al., 2004; Yeo & Friend, 2014; Fasano et al., 2025) and in drop microfluidics (Guttenberg et al., 2004; Brunet et al., 2007, 2009, 2010). It usually dominates contributions to the flow in the bulk fluid arising from the boundary layer type acoustic streaming mechanisms, such as Shlichting or Rayleigh streaming described above. An additional effect related to the interaction of acoustic waves with a surface, the acoustic radiation pressure (Hertz & Mende, 1939; Hamilton & Blackstock, 1998), may be also relevant in the presence of interfaces, such as the fluid-air surface. This effect has been discussed extensively in the literature (King, 1934; Campbell, 1970; Chu & Apfel, 1982; Shiokawa et al., 1989; Hasegawa et al., 2000; Borgnis, 2010; Karlsen et al., 2016; Rajendran et al., 2023), and shown capable of deforming and displacing soft interfaces and of governing the dynamics of films (Biwersi et al., 2000; Alzuaga et al., 2005; Issenmann et al., 2006; Rajendran et al., 2022; Marcos et al., 2025). Altshuler & Manor (2015) consider the problem of a partially wetting water drop atop a MHz frequency propagating vibration in the substrate, in the setup similar to the one considered here, although the reduced viscosity and increased surface tension lead to reduced viscous dissipation of the leaky SAW in the fluid, and a necessity to include disjoining pressure to model liquid-solid interactions. Horesh et al. (2019) further consider a vertical setup where gravity becomes important and show that acoustic radiation pressure is negligible for the partially wetting water due to the curvature of the meniscus, which does not support acoustical resonance.
@@
While the literature discussing the interaction of acoustic waves with fluids is clearly extensive, in what follows, we discuss briefly just a few works that are particularly relevant in the present context, in which a SAW traveling in a solid substrate drives flow in a neighboring fluid. In 1970 Campbell & Jones (1970) simulated a SAW in a piezoelectric solid in contact with an ideal (inviscid) fluid. Then, in the late 1980s, Shiokawa et al. (1989) used the ansatz of a harmonic acoustic waveform in the solid based on ideas developed by Nyborg (1952), converted the approach adopted by Campbell & Jones (1970) to analytic expressions, and calculated the resulting acoustic forcing in the adjacent inviscid fluid due to attenuation of SAW in the solid. More recently, Vanneste & Bühler (2011) revisited the work of Campbell & Jones (1970) in the context of viscous fluids and calculated numerically the corresponding acoustic forcing in the fluid.
@@
In the present paper, we focus on developing a model that can be used to formulate a long wave theory applicable to the free surface evolution of thin viscous films and drops due to the presence of a SAW propagating in an underlying solid substrate. To facilitate such a development, we simplify the previously considered approaches in order to isolate the dominant physical effects governing the dynamics of the film. In particular, we simplify the approach used by Campbell & Jones (1970) by assuming damped harmonic waves in the solid. Furthermore, we compare the previous approaches by Vanneste & Bühler (2011) and Shiokawa et al. (1989) to quantify the importance of viscous effects, which were included by the former authors but ignored by the latter. We follow earlier attempts (Brunet et al., 2010) and focus on the contribution of acoustic streaming to the drop dynamics. In principle, one could also consider including the effects of the acoustic radiation pressure at the free surface, as was done in the recent work of Marcos et al. (2025), who used a Monte-Carlo type approach to study dynamics of oil-water films under acoustic forcing. However, in the present case, the main contribution to drop displacement is the acoustic streaming, which efficiently transfers power from the SAW to the translational motion of the drop; hence, we do not include acoustic radiation pressure here. This point will be discussed in further detail in Sec. 3.
@@
The rest of this paper is structured as follows. As a motivation, in Sec. 2 we discuss simple physical experiments that demonstrate the well-known phenomenon of thick oil films and drops displacing under the action of MHz-frequency Rayleigh-type SAWs. These experiments illustrate the process of dynamic wetting and drop displacement powered by bulk acoustic streaming. We then derive the theoretical model that can explain such observations in Sec. 3. Our theoretical results are presented in Sec. 4: we discuss computational results in Sec. 4.1, before analyzing a simplified traveling–wave type model in Sec. 4.2. Sections 3 and 4, together with the appendices, constitute the main novel contributions of our work. Section 5 is devoted to the summary and conclusions. The appendices include (A) details of the theoretical results, (B) various simplifications and approximations including the inviscid limit, and (C) the outline of the numerical approach implemented. Supplementary materials provide videos of selected experiments and animations of representative computational results.
@@ 2 Experiment
The experiments described here are illustrative and portray the well-known phenomenon in which MHz-frequency surface acoustic waves (SAWs) are used to displace drops and thick films of liquid along the path of the SAW. The purpose of these experiments is to demonstrate the system under investigation, highlight key features and trends, and provide results that will be used to motivate the theory and simulations presented in the following sections.
[…]
@@
In our experiments, we represent the SAW strength (intensity) by measuring the normal displacement amplitude, $A_{\rm n}$, at the solid surface before any attenuation in the liquid, so that $A_{\rm n}$ is the amplitude of the transverse component of the SAW in the substrate. We measure $A_{\rm n}$ over a surface of $1\times 1~\text{mm}^{2}$ about $2$ mm away from the SAW actuator using a scanning laser Doppler vibrometer (MSA-500, Polytech). We control this amplitude by changing the applied voltage, $V$, of the signal generator at the surface of the actuator: consistent with the literature (Ballantine et al., 1996), there is a linear relationship between $A_{\rm n}$ and $V$; see Li et al. (2025) for more details.
[…]
@@
Figure 2: Two sets of images from experiments monitoring silicone oil drops powered by SAWs of three different measured amplitudes $A_{\rm n}$ (1.2 nm–1.4 nm), where (A-C) silicone oil starts from the same location, moves in the direction of the SAW, and reaches different distances (a-c) after 4.09 seconds. The corresponding side-view videos at different acoustic power levels are used to analyze the speed and profile of silicone oil during the movement. (D, d) show top view snapshots (corresponding to (C, c)) of the oil. A thin film of oil, hardly visible in the side view, can be seen more clearly behind the main body of silicone oil in the top views.
[…]
@@
In the experiments with SAW actuation we employ the side view video, recorded by a goniometer camera, to capture the position of the advancing three-phase contact line of the SAW-actuated oil film using the open-source software Tracker (Brown & Cox, 2009) (see Fig. 3(a) for typical data). We also measure the maximum height of the moving oil film using Tracker (typical results in Fig. 3(b)). Figure 3(a) shows that, as expected, the larger the SAW amplitude $A_{\rm n}$, the faster the oil film moves due to the stronger induced leakage wave (Shiokawa et al., 1989). Another notable aspect of SAW-induced dynamics is the contact line speed for early times (<1<1 s): initially, it is rather small, but later increases and then remains constant. This change of contact line speed appears to be a consequence of the initial change of shape of the drop (see also Fig. 1); while this is taking place, contact line speed is small. After the initial deformation, the drop shape remains unchanged, translating approximately uniformly. Figure 3(b) shows consistent results for the maximum drop height, with an increase in height for early times, and constant values for longer times. Note that this uniform translation and constant film height differ significantly from the observed behavior for spreading under a body force such as gravity. These experimental observations will be explained and discussed within the framework of our theoretical model, which is described in the next section. We note that in our experiments, we do not observe drop oscillations as in the experiments using water drops (Brunet et al., 2010), possibly due to the much higher viscosity of the silicone oil.
[…]
@@ 3 Model
We consider a two-dimensional Newtonian fluid layer occupying a domain in the (x,z)(x,z)–plane, traversed by a sound wave generated by a Rayleigh-type-SAW traveling in a neighboring solid; see Fig. 5 for a sketch.
Figure 5: Schematic of a liquid drop (blue curve) driven by a leaky SAW of amplitude $A_{\rm n}$. The SAW travels from left to right in the solid substrate. The drop extends from the rear to the front contact lines at $x_{\rm r}(t)$ and $x_{\rm f}(t)$, respectively. The thin red line represents the amplitude $A_{\rm n}$ of the SAW at the substrate and shows how it is attenuated due to the presence of the liquid (the black dotted line is the envelope of the decaying amplitude in the liquid). The Rayleigh angle is denoted by $\theta_{R}$. The thickness $h^{\ast}$ and the positions $x^{\ast}_{1,2}$ are defined later in the text, see Eq. (35).
We assume that the associated acoustic field Mach number, $M_{a}\equiv\omega A_{\rm n}/c$ (where $\omega$, $A_{\rm n}$ and $c$ are the SAW angular frequency, characteristic normal displacement amplitude at the solid surface, and the phase velocity of the acoustic field in the fluid, respectively) is small ($M_{a}\ll 1$), and that the velocity ($\mathbf{v}$), pressure ($p$), and density ($\rho$) fields in the fluid can be written as
@@
where the leading order components of these fields are associated with quiescent liquid ($\mathbf{v}_{0}=0$, $p_{0}$ constant) of constant ambient density $\rho_{0}$; the first-order corrections ($\mathbf{v}_{1},~p_{1},~\rho_{1}$), represent an oscillatory flow associated with the acoustic field generated by the SAW that average to zero over the period of the acoustic forcing; and the second-order corrections ($\mathbf{v}_{2},~p_{2},~\rho_{2}$) represent the induced steady streaming flow. The assumption implicit in the above expansions in the Mach number is that the barred quantities are comparable in size to the corresponding leading-order contributions. The problem is governed by the Navier–Stokes (NS) and continuity equations,
@@
where $t$ is time; $\mu$, $\mu_{b}$ are the shear and dilatational coefficients of the bulk viscosity, respectively; and $\mathbf{g}=-g\mathbf{e}_{z}$ is the gravitational acceleration. In terms of the above expansions, we will assume that the gravitational term does not appear until second-order in $M_{a}$, which amounts to an assumption that $\mu v_{1}/\ell^{2}\gg\rho g$ (where $v_{1}$ is a representative value of $|\mathbf{v}_{1}|$ and $\ell$ is a representative droplet size), ensuring that gravity is negligible in both leading-order and first-order problems. This condition was checked a posteriori and found to hold. With these assumptions, we substitute the expansions from Eq. (3) into Eq. (2) and collect terms of like orders in $M_{a}$; Secs. 3.1 and 3.2 discuss the first and second order terms, respectively.